OpenStudy (anonymous):

Gary is buying a $1,250 computer on the installment plan. He makes a down payment of $150. He has to make a monthly payment of $48.25 for 2 1/2 years. What is the finance charge?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Exactly the same method as the previous one.

OpenStudy (anonymous):

so 1250 * 55?

OpenStudy (anonymous):

+ 150?

OpenStudy (aum):

Monthly payment = $48.25 Duration of payment = 2 1/2 years = 2.5 * 12 months = 30 months Total monthly payment = ? Add down payment of $150 to above. That will be the total amount paid. Subtract $1,250 to find the finance charge.